x 1 2 3 4 5 6
y 1007 1799 2872 5027 8022 13619

Use linear regression to find the equation for the linear function that best fits this data. Round to two decimal places.

y^( y-hat) = ?

Hint: Enter the data into L1 And L2 then in STAT CALC select 8:LinReg(a+bx)
Answer #1

Sum of X = 21
Sum of Y = 32346
Mean X = 3.5
Mean Y = 5391
Sum of squares (SSX) = 17.5
Sum of products (SP) = 41942

Regression Equation = ŷ = bX + a

b = SP/SSX = 41942/17.5 = 2396.69

a = MY - bMX = 5391 - (2396.69*3.5) = -2997.4

ŷ = 2396.69X - 2997.4